OpenStudy (anonymous):

I would like to create a rectangular vegetable patch. The fencing for the east and west sides costs $4 per foot, and the fencing for the north and south sides costs only $2 per foot. I have a budget of $128 for the project. What are the dimensions of the vegetable patch with the largest area I can enclose?

jimthompson5910 (jim_thompson5910):

let x = length and y = height, so we might have something like this 

jimthompson5910 (jim_thompson5910):

The fencing for the east and west sides costs $4 per foot so it costs 4y + 4y = 8y dollars to do the east and west sides

jimthompson5910 (jim_thompson5910):

the fencing for the north and south sides costs only $2 per foot so it costs 2x+2x = 4x dollars to do the north and south sides

jimthompson5910 (jim_thompson5910):

In total, it will cost you 4x + 8y dollars to do the whole enclosure

jimthompson5910 (jim_thompson5910):

Your budget is $128, so let's say you spend exactly that much money (nothing more, nothing less)

jimthompson5910 (jim_thompson5910):

this would mean 4x + 8y = 128

jimthompson5910 (jim_thompson5910):

Solve for x or y and tell me what you get

jimthompson5910 (jim_thompson5910):

it might make sense to solve for y, but either one works

OpenStudy (anonymous):

like this? y=16-.5x

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

keep that in mind because we'll use that later

jimthompson5910 (jim_thompson5910):

now the area of this rectangle  is A = Length * Width A = LW A = x*y